Two long parallel wires are 10 cm apart and carry currents of 6 A and 4 A. what is the field B at 4 A carrying wire due to 6 A wire.
1
Expert's answer
2020-05-14T09:02:19-0400

The magnetic field created by a long wire with current

"B=\frac{\mu_0 I}{2\pi r}\\
=\frac{4\pi\times 10^{-7}\times 6}{2\pi \times 0.1}=1.2\times 10^{-5} \:\rm T"